# A New Approach to the Representation Theory of Lorentzian Pseudo-Tensors

### v1

@@ Abstract
A novel approach to the finite dimensional representation theory of the entire Lorentz group $\operatorname{O}(1,3)$ is presented. It is shown that the entire Lorentz group may be understood as a semi-direct product between the identity component of the entire Lorentz group, and the Klein four group of reflections: $\operatorname{O}(1,3)=\operatorname{SO}^{+}(1,3)\rtimes\operatorname{K}_{4}$. The discussion concludes with the convenient representation theory of generic tensor representations of $\operatorname{O}(1,3)$, namely that there are four physically meaningful representations of $\operatorname{O}(1,3)$ for each representation of $\operatorname{SO}^{+}(1,3)$. There is a brief discussion of the time reversal of the electromagnetic field, concluding in agreement with standard texts such as Jackson JACKSON3RD , and works by Malament MALAMENT2004 .
I What is meant by ‘Time Reversal’?
@@
Time reversal is one of the least well understood topics in modern physics, with students and experts alike often having little understanding of the intricacies involved in defining the notion. Most sources will merely point back to the progenitor of the understanding of time reversal, Wigner’s exploration of the topic in the early 1930’s Wigner1932 , WIGNERBOOK . While certain notions have been set in stone since (such as time reversal’s necessary anti-unitarity), there is still great deal of disagreement about particular notions, interpretations, and applications of the concept. Furthermore despite physics textbooks being largely in agreement about the effects a ‘time reversal operator’ should have on particular physical phenomena, there remains no universally agreed upon definition for what physicists and philosophers of physics mean by when they say the ‘time reversal’ of a system. This is no better highlighted than in the disagreements and debate between, for example, Malament MALAMENT2004 , Arntzenius & Greaves GREAVESARNTZENIUS , Roberts ROBERTS2022 , and Albert ALBERT2003 , and Callender Callender .
@@
This confusion around the topic led the author to try and understand the discrete spacetime symmetries in a self consistent way, and this paper is the result of that investigation. What is meant mathematically in this paper by ‘time reversal’ will be made clear in the next section. In words, it should be understood literally as a reflection, directly analogously with parity. To give a simple explicit example, suppose we have a certain ordering of objects in space, three spheres identical but in color, named $r,g,b$, placed at coordinates $(1,2,3)$ along an axis directly in front of us. Placing a mirror perpendicular to the axis, at coordinate $(4)$ will provide us an image of the spheres, under the operation of parity, the images appearing at coordinates $(5,6,7)$ with colours $b,g,r$ respectively. What we mean by a system being parity symmetric, is that were we to understand the parity reflected system as a system within in our world, the two systems appear identical, or essentially identical up to the application of other agreed upon transforms. Importantly, note that when we consider the parity reflected state, we do not parity transform ourselves as well. Doing so would result in the trivial fact that all states appeared to us as parity symmetric. In the example given, the parity reflected state is not strictly symmetric, unless we mean symmetric up to translations and rotations, since the order of the colours has been reversed, and the images coordinates are different than the originals. However were we to replace the third sphere $b$ with another sphere $r$, then the state is manifestly symmetric (or invariant) under a reflection through coordinate $2$: i.e. the parity reflection of this setup, understood as another possible setup of the objects in our world, is identical to the initial setup of the objects.
@@
The philosophy is the same when it comes to temporal reflection. The one dimensional nature of time necessitates a partial ordering of events for any observer, and so under a temporal reflection, what is to be inspected for symmetry or invariance, is the set of events in reversed order, but understood as a possible sequence flowing forward in time. To be explicit here let it be said we do not interpret a time reversal of an ordered set of events, as the events evolving backward in time. We interpret it rather as a possible forward time evolution of some (possibly differently prepared) system.
@@
Finally in the work below, what is meant by ‘time reversal invariance’, ‘parity symmetry’, and synonyms thereof, is not the strict examples provided above: it is not that states or systems are essentially identical before and after the application of the reflection. What is more interesting physically and almost always what is meant by these terms, is that given a physical theory describing the evolution of a system: a theory exhibits parity symmetry if for every solution to the theory’s equations of motion, any parity reflection of the solution is also a solution to the equations of motion. Directly analogously a theory is said to exhibit time reversal invariance if for every solution to the theory’s equations of motion, any temporally reflected solution is a solution to the equations of motion. In less precise language, for a system described by some physical theory, if all predicted behaviours of the system as seen in a mirror are also possible behaviours of the system, then the theory is parity symmetric. Likewise if for all predicted behaviours of the system, when the order of events is reversed, these temporally reflected behaviours are also possible behaviours of the system, the theory is time reversal symmetric. [^1] The author presently only has classical and semi-classical pictures in mind. The full quantum mechanical case requires additional care and detail, for several reasons, foremost among them being that the discussion becomes about operators, state spaces, and observables, instead of trajectories and values of fields.
II Representations of Parity & Time Reversal Operators
In the following I propose that the parity and time reversal properties of the objects which enter our theories, may be understood merely from the representation theory of the entire Lorentz group: $\operatorname{O}(1,3)$. The word entire here is used to distinguish from the usual abuse of language whereby physicists say ‘The Lorentz Group’ to mean only the component of the entire Lorentz group continuously connected to the identity transform. The entire group has four connected components, disjoint from one another, illustrated in Fig. (1).
Figure 1: The four connected components of the group $\operatorname{O}(1,3)$, along with the discrete reflections mapping between them. In the defining representation $\mathcal{R}$ = $\operatorname{diag}\{1,-1,-1,-1\}$.
We avoid the usual identification of the parity operator $\mathcal{P}$ with $\mathcal{R}$, and the time reversal operator $\mathcal{T}$ with $-\mathcal{R}$. The reason for this avoidance is precisely because the representation theory of the entire Lorentz group is more subtle than this simple assignment of $\mathcal{P}$ and $\mathcal{T}$: we will come to see these operators behave differently when acting upon distinct representations, just as say, familiar angular momenta operators look different acting upon different representations. The entire Lorentz group group has $3$ important proper subgroups, seen in Fig (2).
$\mathcal{R}$$\mathcal{R}$
111111-1- $-1$
$1$
Figure 2: Three proper Lie subgroups of $\operatorname{O}(1,3)$. They are in order: The projective orthogonal group $\operatorname{PO}(1,3)$, the special orthogonal group $\operatorname{SO}(1,3)$, and the projective special orthogonal group $\operatorname{PSO}(1,3)$, also known as the proper orthochronus Lorentz group.
The projective orthogonal group, $\operatorname{PO}(1,3)$, should be understood as the ‘centerless’ [^2] The center of a group is the set of all elements of the group which commute with every other element in the group AbsAlg . For $\operatorname{O}(1,3)$ the center is {$\{\mathcal{I},-\mathcal{I}\}$. orthogonal group: in this subgroup the action of two Lorentz transforms related by multiplication of $-1$, is indistinguishable. This is why we are missing the would be ‘fourth’ subgroup: in this subgroup $\mathcal{R}$ and $-\mathcal{R}$ act identically. An example of an object which transforms under this subgroup is four-momentum, since time reversal and parity inversion have the same effect upon the vector. The special orthogonal group $\operatorname{SO}(1,3)$ should be more familiar: it is all those Lorentz transforms with determinant of $+1$. Of note is that due to the non-compactness of the group, we still find two components of this sub group: those transforms which (in the defining representation) preserve the orientation of time and space separately, and those which reverse both simultaneously.
The intersection of these two subgroups leads to the projective special orthogonal group, known to its friends as the proper orthochronus Lorentz group:
$\operatorname{SO}^{+}(1,3)\equiv\operatorname{PSO}(1,3).$ (1)
This is the piece most physicists mean when they say ‘the Lorentz group’, it is the identity component of the group $\operatorname{O}(1,3)$: it is the piece of the Lorentz group we acquire by exponentiation of its Lie algebra, and so it is the piece for which the standard representation theory is well understood. What we are missing then is the representation theory of the three other components of the entire Lorentz group; but these components are merely ‘reflected copies’ of the identity component. Therefore what is desired is a way to think of the entire Lorentz group as some product of the piece we understand, and some discrete reflection group which generates these reflected copies. In what is to follow we will come to understand the entire Lorentz group may be understood as the semi-direct product of groups [^3] See Appendix (A) for a review of semi-direct products.
Where $\operatorname{K}_{4}$ is the so called Klein four group vier , corresponding to the action of the discrete reflections $\mathcal{P},\mathcal{T}$, and $\mathcal{PT}$, explored in detail in the next section. We will often refer to this group as “the discrete reflection group”. What is particularly important about the above decomposition of the entire Lorentz group, is that the representations of a products of groups may often be given by the tensor product of representations of said groups. Therefore to understand representations of the entire Lorentz group, we must understand the representations of $\operatorname{K}_{4}$. The physical distinction between say, a pseudo-four-vector and a polar four-vector, can only be captured by understanding the representation theory of the entire Lorentz group. Failing to do this leads to misunderstandings about the nature of objects under the operation of $\mathcal{P}$ and $\mathcal{T}$.
II.1 Simplest Motivating Example: Velocities
@@
It is illustrative to see how velocities naturally lead to a consideration of something like ‘charges’ under the discrete reflection group. Given a one dimensional trajectory $x(t)$, and its velocity function $u(t)=dx/dt$, a temporal reflection about $t=0$ gives us the corresponding function $x(-t)$. If we were naively given the function $u(t)$ and asked to perform a time reflection of it, we would just as well return the function $u(-t)$. However, this function is not the velocity of the time reversed trajectory:
@@
Of course we can see chain rule from the time derivative is what causes the additional sign flip, but often we find ourselves in situations where this ‘implicit’ additional time dependence is not obvious. If we wish to understand the time reversal of some complicated dynamics, it is cumbersome to unpack the definition of each object, to see if a possible chain rule might apply. Furthermore, physically it is not only those objects involving spacial derivatives which pick up a sign under spacial reflections. As such we should anticipate the possibility of objects which pick up a sign under temporal reflections whether or not their definition explicitly contains a time derivative. This motivates the notion that some objects are ‘charged’ or ‘odd’ under time reversal, and some objects are ‘uncharged’ or ‘even’. Keeping this notion around and understanding its group structure allows for simple book keeping of objects which physically ought to pick up additional signs under reflections generally. In this case positions would be uncharged under time reversal and velocities would be charged, whereas they are both charged under parity inversion.
II.2 The Klein Four group
@@
The Klein four group $\operatorname{K}_{4}$, also known as the Vierergruppe, abstractly is a discrete Abelian group containing four elements, suggestively named $\{1,\mathcal{P},\mathcal{T},\mathcal{PT}\}$ satisfying:
@@
As $\operatorname{K}_{4}$ is Abelian, Schur’s lemma WuKiTung tells us its irreducible representations over $\mathbb{C}$ are all one dimensional. For $\operatorname{K}_{4}$ this remains true even for real representations. There are exactly four distinct irreducible representations of the Klein group, given in Tab. (1).
@@
For convenience let us introduce a charge operator $K$ for the Klein group: when given an object, $K$ returns the representation of the Klein group the object transforms under, i.e. its charges under the discrete reflection group. In the following section we will consider larger representations of $\operatorname{K}_{4}$, which from the above know must take the form of direct sums of one dimensional representations. From here on, we will refer to representations simply by their odd charges, i.e. their subscript $1,P,T,PT$, and the abstract operators (the things being represented) will be written $\mathcal{P}$ and $\mathcal{T}$.
II.3 The Defining Representation of $\operatorname{O}(1,3)$
The defining representation of $\operatorname{O}(1,3)$ acting on $\mathbb{R}^{1,3}$ with coordinates $x^{\alpha}=(t,\vec{x})$ identifies the parity operator $\mathcal{P}$, and the time reversal operator $\mathcal{T}$ with:
@@ II.3 The Defining Representation of $\operatorname{O}(1,3)$
From the knowledge above, we know these operators must be the direct sums of irreducible one dimensional representations. Inspection shows us these are the action of $\mathcal{P}$ and $\mathcal{T}$ upon the representation $T\oplus P\oplus P\oplus P=T\oplus 3P$. That is to say the representation which standard Cartesian coordinate $4$-vectors take under $\operatorname{K}_{4}$ is given by one component which is odd under time reflections (the time coordinate), and three components which are odd under parity (the spacial coordinates), i.e.:
$K[x^{\alpha}]=\begin{pmatrix}T\\ P\\ P\\ P\end{pmatrix}:=\begin{pmatrix}T\\ 3P\end{pmatrix}:=\begin{pmatrix}T\\ \vec{P}\end{pmatrix}.$ (6)
Hopefully the abuses of notation are clear. This should all be rather straightforward and intuitive at this point. Let us inspect something more interesting such as the momentum of a point particle. Suppose a point particle has a trajectory given in coordinates by
@@
$\begin{split}K[p^{\alpha}]&=K[\mu]\otimes K\left[\frac{d}{d\tau}\right]\otimes K[X^{\alpha}]\\ &=1\cdot T\cdot\begin{pmatrix}T\\ \vec{P}\end{pmatrix}=\begin{pmatrix}1\\ \vec{PT}\end{pmatrix}.\end{split}$ (10)
What this tells us is that the energy is completely invariant under the action of the discrete reflection group $\operatorname{K}_{4}$, and the momenta are odd under both parity and time reflections. This matches our intuition, and is the first example of a failure of the ‘standard’ defining assignments of the parity and time reversal operators. In this representation both $\mathcal{P}$ and $\mathcal{T}$ are represented by the same matrix, ‘$\mathcal{R}$’ (defined above), and $\mathcal{PT}$ acts as the identity. From a representation theoretic point of view momentum vectors may be thought of as taking representations of $\operatorname{PO}(1,3)$, instead of $\operatorname{O}(1,3)$.
There are still unanswered questions at this point, such as how the discrete reflection group interacts with the (proper orthochronus) Lorentz group, especially under the action of boosts, which at present would seem to ‘mix up’ our clean one dimensional decomposition of the four vectors, possibly leading to different observers ascribing different charges under the reflection group to the same physical objects. In the next section we will explicitly construct the semi-direct product $\operatorname{O}(1,3)=\operatorname{SO}^{+}_{\uparrow}(1,3)\rtimes\operatorname{K}_{4}$, and address these concerns.
II.4 The Effect of Parity and Time Reversal on Boosts and Rotations
If in one frame we have a particle traveling along the $x$ axis with velocity $\beta=v/c$, we may transform our coordinates to a co-moving frame by application of the boost $B(\beta)=e^{\omega K_{x}}$, where $\omega=\operatorname{arctanh}(\beta)$ is the rapidity, and $K_{x}$ is the relevant generator of boosts defined in Appendix (B). Under either a parity inversion or time reflection of this setup, the velocity will be reversed, [^5] Note that $\operatorname{arctanh}$ is an odd function. and so the correct boost to enter the co-moving frame after either reflection will be given via $B(-\beta)=e^{-\omega K_{x}}$. Of course, if one boosts first, and then reflects, it is not different than reflecting first, and boosting in the reversed direction after. This can be phrased via commutation relations:

### v2

# A Convenient Representation Theory of Lorentzian Pseudo-Tensors:

@@ Abstract
A novel approach to the finite dimensional representation theory of the entire Lorentz group $\operatorname{O}(1,3)$ is presented. It is shown how the entire Lorentz group may be understood as a semi-direct product between its identity component and the Klein four group of spacetime reflections: $\operatorname{O}(1,3)=\operatorname{SO}^{+}(1,3)\rtimes\operatorname{K}_{4}$. This gives way to a convenient classification of tensors transforming under $\operatorname{O}(1,3)$, namely that there are four representations of $\operatorname{O}(1,3)$ for each representation of $\operatorname{SO}^{+}(1,3)$, and it is shown how the representation theory of the Klein group $\operatorname{K}_{4}$ allows for simple book keeping of the spacetime reflection properties of general Lorentzian tensors, and combinations thereof, with several examples given. There is a brief discussion of the time reversal of the electromagnetic field, concluding in agreement with standard texts such as Jackson JACKSON3RD , and works by Malament MALAMENT2004 .
I What do we mean by ‘Time Reversal’?
@@
Time reversal is one of the most confused topics in modern physics, with students and experts alike often having little understanding of the intricacies involved in defining the notion, so it is worth specifying precisely what is meant by it. Most sources will simply point to the progenitor of the modern understanding of time reversal, Wigner’s exploration of the topic in the 1930’s Wigner1932 , WIGNERBOOK . While certain notions have been set in stone since (such as time reversal’s necessary anti-unitarity), there remains a great deal of confusion about particular interpretations and applications of the concept. Despite physics textbooks being largely in agreement about the effects a ‘time reversal operator’ should have on any particular physical phenomenon, it is often treated case-by-case, and there remains no universally agreed upon definition for what physicists and philosophers of physics mean when they discuss the ‘time reversal’ of a system. This is no better highlighted than in the disagreements and debate between, for example, Malament MALAMENT2004 , Arntzenius & Greaves GREAVESARNTZENIUS , Roberts ROBERTS2022 , Albert ALBERT2003 , and Callender Callender . It is this confusion which led me to attempt to understand the discrete spacetime symmetries in a self consistent way. I believe the attempt to have been successful, making precise many details lacking in textbook explanations, and delivering a convenient formalism for keeping track of the properties of physical quantities under spacetime reflections.
@@
What is meant mathematically in this paper by ‘time reversal’ will be made clear in Sec. (II). In words however, it should be understood literally as a reflection directly analogous to parity. To give a simple example, suppose we have an ordering of objects in space, three spheres identical but in color named $r,g,b$, placed at coordinates $(1,2,3)$ along an axis directly in front of us. Placing a mirror normal to the axis at coordinate $(4)$ produces an image of the spheres under the operation of parity, the images appearing at coordinates $(5,6,7)$ with colors $b,g,r$ respectively. When we talk about a system, object, theory, etc. being parity symmetric, we mean that were we to understand the parity reflected system as if it were a system within our world, the two systems appear identical, or essentially identical up to the application of other agreed upon transforms and symmetries. In the example given the parity reflected state is not strictly symmetric, unless we mean symmetric up to translations and rotations, since the order of the colors has been reversed, and the images coordinates are different than the originals. Were we to replace the third sphere $b$ with another sphere $r$, then the state is manifestly symmetric (or invariant) under a reflection through coordinate $2$: i.e. the parity reflection of this setup, understood as another possible setup of the objects in our world, is identical to the initial setup of the objects. Importantly, note that when we consider the parity reflected state, we do not parity transform ourselves as well.
@@
The philosophy is the same when it comes to temporal reflection. The one dimensional nature of time necessitates a partial ordering of events for any observer, and so under a temporal reflection, what is to be inspected for symmetry or invariance, is the set of events in reversed order, but understood as a possible sequence of events flowing forward in time. To be explicit here let it be said we do not interpret a time reversal of an ordered set of events, as the events evolving backward in time. We interpret it rather as a possible forward time evolution of some (possibly differently prepared) system.
@@
Finally in the work below, what is meant by ‘time reversal invariance’, ‘parity symmetry’, and synonyms thereof, are not the strict examples provided above: it is not that states or systems are essentially identical before and after the application of the reflection. What is more interesting physically and typically what is meant by these terms, is that given a physical theory describing the evolution of a system: a theory exhibits parity symmetry if for every solution to the theory’s equations of motion, any parity reflection of the solution is also a solution to the equations of motion. Directly analogously a theory is said to exhibit time reversal invariance if for every solution to the theory’s equations of motion, any temporally reflected solution is a solution to the equations of motion.
In less precise language, for a system described by some physical theory, if all predicted behaviors of the system as seen in a (any) mirror are also possible behaviors of the system, then the theory is parity symmetric. Likewise if for all predicted behaviors of the system, when the order of events is reversed these temporally reflected behaviors are also possible behaviors of the system, the theory is time reversal symmetric. [^1] The author presently only has classical and semi-classical pictures in mind. The full quantum mechanical case requires additional care and detail, for several reasons, foremost among them being that the discussion becomes about operators, state spaces, and observables, instead of trajectories and values of fields. Maxwell’s equations, for example, exhibit both of these symmetries; a proof of this in given in appendix. (A.4).
I.1 Simplest Motivating Example: Velocities
@@
Before we come to the formalities of group and representation theory in the following sections, it is illustrative to see how something as simple as velocities lead naturally to a consideration of something like ‘parity’ or ‘charges’ under reflections of coordinates. Given a one dimensional trajectory $x(t)$, and its velocity function $u(t)=dx/dt$, a temporal reflection about $t=0$ gives us the corresponding function $x(-t)$. If we were naively given the function $u(t)$ and asked to perform a time reflection of it, we would just as well return the function $u(-t)$. However, this function is not the velocity of the time reversed trajectory:
@@
Of course we can see the chain rule from the time derivative is what causes the additional sign flip, but often we find ourselves in situations where this ‘implicit’ additional time dependence is not obvious. Should we wish to understand the time reversal of some complicated dynamics, it is cumbersome to unpack the definition of each object, to see if a possible chain rule might apply. Integrations and and derived quantities thereof generally should invite us to anticipate the possibility of objects which pick up a sign under temporal reflections whether or not their definition explicitly contains a time coordinate or time derivative. This motivates the notion that some objects are ‘charged’ or ‘odd’ under time reversal, and some objects are ‘uncharged’ or ‘even’; no different from how the parity of a vector distinguishes coordinate vectors and pseudo-vectors. We will see that treating this notion rigorously allows for simple book keeping of quantities which physically ought to pick up additional signs under spacetime reflections. In this language, under time reversal positions would be uncharged (even) and velocities would be charged (odd), whereas they are both charged (odd) under spacial inversion.
II The Structure of O(1,3)
Figure 2: The four connected components of the group $\operatorname{O}(1,3)$, along with the reflections mapping between these components: $\mathcal{R}$ = $\operatorname{diag}\{1,-1,-1,-1\}$.
In the following sections a convenient representation theory of Lorentzian pseudo-tensors [^2] A pseudo-tensor is used here in the sense as pseudo-vector and pseudo-scalar: additional sign changes under the action of reflections. Axial and pseudo are synonymous prefixes. is constructed: a way of labeling physical quantities so their the parity and time reversal properties are transparent to us at a glance. For brevity it is assumed the reader is familiar with special relativity, in particular the ideas of inertial observers, a lack of preferred reference frames, the notion of Lorentz invariance, and boosts and rotations as the building blocks of Lorentz transforms.
This approach comes out of the representation theory of the entire Lorentz group: $\operatorname{O}(1,3)$. The word entire here is used to distinguish from the usual abuse of language whereby physicists say ‘The Lorentz Group’ to mean only the component of the entire Lorentz group connected to the identity transform.
$\operatorname{SO}^{+}(1,3)\equiv\operatorname{PSO}(1,3).$ (2)
The identity component is known to its friends as the proper orthochronus Lorentz group, a mathematician would call it the projective special orthogonal group. The entire group has four connected components, disjoint from one another, illustrated in Fig. (2). Note that the usual identifications of the parity operator $\mathcal{P}$ with $\mathcal{R}$ and the time reversal operator $\mathcal{T}$ with $-\mathcal{R}$, are avoided. The reason for this avoidance is precisely because the representation theory of the entire Lorentz group is more subtle than this simple assignment of $\mathcal{P}$ and $\mathcal{T}$; we will come to see these operators behave differently when acting upon distinct representations, just as say, familiar angular momenta operators look different acting upon different representations.
Any Lorentz transform in $\operatorname{SO}^{+}(1,3)$, the identity component, can be reached by appropriate choice of angle and boost parameter. [^3] Formally this is declaring the fact that the map from the Lie algebra to the identity component of the Lie group is surjective. It is true here but not always the case, such as for $\operatorname{SL}(2,\mathbb{C})$ These transforms preserve orientations in space and the sense of time. Elements within the other three disconnected components may then be reached by multiplication of one of the reflections: $\{\mathcal{R},-\mathcal{R},-1\}$, which allow for changes of frame which swap the handedness of a coordinate system, or reverse the sense of time. These three reflections together with the identity form a group known as the Klein group $\operatorname{K}_{4}$, discussed next in Sec. (II.1).
The identity component is the piece of the entire group for which the standard representation theory is well understood. What we are missing then is a compelling representation theory of the three other components of the entire Lorentz group; but these components are merely ‘reflected copies’ of the identity component. Therefore what is desired is a way to think of the entire Lorentz group as some product of the piece we understand, and some discrete reflection group which generates these reflected copies. In what is to follow we will see the entire Lorentz group may be understood as the semi-direct product of groups [^4] See Appendix (B.1) for a review of semi-direct products.
Again $\operatorname{K}_{4}$ is the Klein four group vier , corresponding to the action of the discrete reflection operators $\mathcal{P},\mathcal{T}$, and $\mathcal{PT}$, explored in detail in the next section. We will often refer to this group as “the discrete reflection group”. What is particularly important about the above decomposition of the entire Lorentz group, is that the representations of a products of groups can often be given by the tensor product of representations of said groups, and this is true here. Therefore to understand representations of the entire Lorentz group, we must understand the representations of $\operatorname{K}_{4}$. The physical distinction between say, a pseudo-four-vector and a coordinate four-vector, can only be captured by understanding the representation theory of the entire Lorentz group, and failing to do this can lead to misunderstandings about the nature of objects under the operation of $\mathcal{P}$ and $\mathcal{T}$.
II.1 The Klein Four Group
@@
The Klein four group $\operatorname{K}_{4}$, also called the Vierergruppe, abstractly is a finite Abelian group containing four elements, suggestively named $\{1,\mathcal{P},\mathcal{T},\mathcal{PT}\}$ satisfying:
@@
As $\operatorname{K}_{4}$ is Abelian, Schur’s lemma WuKiTung tells us its irreducible representations over $\mathbb{C}$ are all one dimensional. For $\operatorname{K}_{4}$ this remains true even over $\mathbb{R}$. There are exactly four distinct irreducible representations of the Klein group, given in Tab. (1).
@@
For convenience let us introduce a charge operator $K$ for the Klein group: when given an object, $K$ returns the representation of the Klein group the object transforms under, i.e. its ‘charges’ or parities under the discrete reflection group. In the following section we will consider larger representations of $\operatorname{K}_{4}$; these must take the form of direct sums of one dimensional representations. From here on, we will refer to representations simply by their odd charges, i.e. their subscript $1,P,T,PT$, and the abstract operators (the things being represented) will be written $\mathcal{P}$ and $\mathcal{T}$.
II.2 The Defining Representation of $\operatorname{O}(1,3)$
The defining representation of $\operatorname{O}(1,3)$ acting on $\mathbb{R}^{1,3}$ with Cartesian coordinates $x^{\alpha}=(t,\mathbf{x})$ identifies the parity operator $\mathcal{P}$, and the time reversal operator $\mathcal{T}$ acting on these coordinates via the matrices $\mathcal{P}_{\operatorname{def}}$ and $\mathcal{T}_{\operatorname{def}}$:
@@ II.2 The Defining Representation of $\operatorname{O}(1,3)$
This representation is given by one component which is odd under time reversal (the time coordinate), and three components which are odd under parity (the spacial coordinates). The reflection charges of the vector are:
$K[x^{\alpha}]=T\oplus P\oplus P\oplus P=\begin{pmatrix}T\\ P\\ P\\ P\end{pmatrix}:=\begin{pmatrix}T\\ \vec{P}\end{pmatrix}.$ (6)
Let us inspect something more interesting: the momentum of a point particle. Suppose a point particle has a trajectory given in coordinates by
@@
$\begin{split}K[p^{\alpha}]&=K[\mu]\otimes K\left[\frac{d}{d\tau}\right]\otimes K[X^{\alpha}]\\ &=1\cdot T\cdot\begin{pmatrix}T\\ \vec{P}\end{pmatrix}=\begin{pmatrix}T^{2}\\ T\cdot\vec{P}\end{pmatrix}=\begin{pmatrix}1\\ \vec{PT}\end{pmatrix}.\end{split}$ (10)
This tells us that the energy is completely invariant under the action of the discrete reflection group $\operatorname{K}_{4}$, and the three-momenta are odd under both parity and time reflections. This matches our intuition, and is the first example of a failure of the ‘standard’ defining assignments of the parity and time reversal operators. In this representation both $\mathcal{P}$ and $\mathcal{T}$ are represented by the same matrix, ‘$\mathcal{R}$’ (defined above), and $\mathcal{PT}$ acts as the identity.
Next we need to make sense of how the discrete reflection group $\operatorname{K}_{4}$ interacts with the proper orthochronus Lorentz group $\operatorname{SO}^{+}(1,3)$, especially under the action of boosts which at present would seem to ‘mix up’ our clean one dimensional decomposition of these four-vectors, perhaps leading to different observers ascribing different charges under the reflection group to the same physical objects and causing a contradiction. The next two sections address this concern, constructing precisely those combinations of reflection charges which have physically reasonable transformation properties.
II.3 The Effects of Parity and Time Reversal on Boosts and Rotations
The interaction of boosts and rotations with the basic spacetime reflections is mostly straightforward. Suppose in one frame we have a particle traveling along the $x$ axis with velocity $\beta=v/c$; we may transform our coordinates to a co-moving frame by application of the boost $B(\beta)=e^{\omega K_{x}}$, where $\omega=\operatorname{arctanh}(\beta)$ is the rapidity, and $K_{x}$ is the relevant generator of boosts defined in Appendix (B.2). Under either a parity inversion or time reversal of this setup, the velocity of the reflected system will be reversed compared to the original, and so the correct boost to enter the co-moving frame after either reflection will be given via $B(-\beta)=e^{-\omega K_{x}}$. Of course, if one boosts first, and then reflects, it is the same as reflecting first and then boosting in the reversed direction. This can be phrased via commutation relations: